OpenStudy (anonymous):

The region R in the first quadrant is enclosed by the lines x=0, y=5 and the graph of y = x^2+1. What is the volume of the solid generated when R is revolved about the y-axis?

OpenStudy (anonymous):

I am trying to use shell method but I seem to not get the right answer. Please help.

OpenStudy (anonymous):

what's your answer?

OpenStudy (anonymous):

My answer was in the 3000's Its nowhere near any of the multiple choice answers. Should I just use disk/washer method instead?

OpenStudy (anonymous):

ok. since you wanted shell method we'll do that...

OpenStudy (anonymous):

my equation I have made so far was 2\[2\pi \int\limits_{1}^{5} [(y)(\sqrt{y}-1)]dy\]

OpenStudy (anonymous):